OpenStudy (anonymous):

How can I prove that the space spanned by two vectors in R^n is a line through the origin, a plane through the origin, or the origin itself?

OpenStudy (turingtest):

i'm very rusty on LA, but let's consider what we need to show that a given set of vectors spans a vector space: that every vector in that space can be written as a linear combination of the given vectors the simplest case is that which only spans the origin. what vectors would only span the origin?

OpenStudy (anonymous):

it would be (0,0)

OpenStudy (anonymous):

or a zero vector

OpenStudy (turingtest):

right, and we can see this by seeing that\[c_1\langle0,0\rangle+c_2\langle0,0\rangle=\vec 0\]now what kind of vectors would span \(\mathbb R^1\) ?

OpenStudy (anonymous):

it would be (0,1) or (1,0)

OpenStudy (turingtest):

well you asked for two vectors, and the don't have to be unit vectors so if we are to choose two vectors that together only span \(\mathbb R^1\) they must have a certain relation... which is what?

OpenStudy (anonymous):

linearly dependent?

OpenStudy (turingtest):

yep :) so we can write this as\[\vec v_1=\langle x,y\rangle\\\vec v_2=\langle c x, cy\rangle\]if you take the determinant of this you will find it's zero, which shows it doesn't span \(\mathbb R^2\)

OpenStudy (turingtest):

it's easy to show that either one individually spans \(\mathbb R^1\) though as for vectors that span \(\vec R^2\) what property must they have?

OpenStudy (turingtest):

\(\mathbb R^2\)*

OpenStudy (anonymous):

it must be closed under addition and scalar multiplication

OpenStudy (turingtest):

well do you know how you would show that for the last case?

OpenStudy (anonymous):

i guess by multiplying it by a scalar, k where k=/= 0?

OpenStudy (anonymous):

@amistre64 any help here?